# 1]ByteDance

### v1

@@ 3.1 Variational Autoencoder
VAE defines the upper bound for generation realism and fidelity. Due to VAE’s function of connecting latent space and pixel space, the reconstruction quality of VAE itself reflects the loss of information by compression and the upper bound of fidelity of generation tasks. As shown in Figures 4 and 5, our VAE model effectively reconstructs fine textures and high-dynamic videos, which may be a primary factor contributing to the high realism and vivid motion in our generated videos.
@@ 3.2 Diffusion Transformer Model
Figure 8 compares the dual-stream [21] and hybrid-stream architectures under the same training compute budget (or equivalently, training FLOPs), where the $y$-axis represents the video diffusion loss and the $x$-axis denotes the compute. The results indicate that the hybrid-stream architecture consistently achieves lower loss compared to the dual-stream architecture. Based on the above designs, we build the 7B hybrid-stream model with a hidden size of 3584 and a total of 32 layers.
@@ 3.3.2 Post-training
After the pre-training phase, we apply supervised fine-tuning (SFT), followed by reinforcement learning from human feedback (RLHF), to further improve the aesthetic quality, motion consistency, and structural coherence of the outputs. The post-training process is carried out independently for the text-to-video and image-to-video tasks.
@@
However, prolonged SFT training can lead to rapid overfitting, resulting in degraded prompt-following ability and reduced motion quality. As illustrated in Figure 14, the model’s ability to follow prompts deteriorates during the SFT training phase.
@@ 4.1 Quantitative Analysis of Video Generation
In the text-to-video task, Seaweed-7B ranks among the top 2-3 models in the Elo comparison. It follows the top-ranked model, Veo 2, performs comparably to Wan 2.1-14B, and outperforms Kling 1.6 (HD). Figure 19 presents a detailed comparison with the two leading models, Veo 2.0 and Wan 2.1. These results demonstrate that the Seaweed model, trained with 665,000 H100 GPU hours, achieves competitive performance compared to larger models trained with significantly greater computational resources.
@@ 7 Conclusion
Below are the researchers (individuals who worked full-time on the project) and contributors (individuals who provided assistance to the project). The Points of Contact (POCs) for each workstream are marked with $\dagger$.
@@
Research Model (alphabetical order): Ceyuan Yang, Fei Xiao, Feng Cheng, Hao Chen, Haoyuan Guo, Meng Wei, Peihao Zhu, Qi Zhao, Shanchuan Lin, Yang Zhao, Zhijie Lin, Zhiwu Qing.
Research Data (alphabetical order): Fangyuan Kong, Feilong Zuo, Jiangqiao Yan, Liangke Gui, Lu Qi, Sen Wang, Sheng Bi, Siyu Zhang, Tuyen Hoang, Xuejiao Zeng, Zhibei Ma, Ziyan Yang.
@@
Contributors: Bingchuan Li, Chao Liang, Deyao Zhu, Gaojie Lin, Gen Li, Hao He, Jianwen Jiang, Jianyi Wang, Jiaqi Yang, Jiawei Liu, Junfei Xiao, Lijie Liu, Lizhen Wang, Longhao Zhang, Qian He, Siyu Zhou, Tianshu Hu, Tianxiang Ma, Weilin Huang, Xiaobin Zhuang, Xiaohui Shen, Xinglong Wu, Yongming Zhu, Yuping Wang, Yuwei Guo, Yuxuan Luo, Yuxuan Wang, Zerong Zheng, Zhengkun Rong, Zhuo Chen, Zhuowei Chen.

### v2

# Seaweed-7B: Cost-Effective Training of Video Generation Foundation Model

@@ 3.1 Variational Autoencoder
Due to VAE’s function of connecting latent space and pixel space, the reconstruction quality of VAE itself reflects the loss of information by compression and significantly affect the fidelity of generation tasks. As shown in Figures 4 and 5, our VAE model effectively reconstructs fine textures and high-dynamic videos, which may be a primary factor contributing to the high realism and vivid motion in our generated videos.
@@ 3.2 Diffusion Transformer Model
Figure 8 compares the dual-stream [21] and hybrid-stream architectures under the same training steps where comparison under the same compute budget is also presented. The results indicate that the hybrid-stream architecture consistently achieves lower loss compared to the dual-stream architecture. Based on the above designs, we build the 7B hybrid-stream model with a hidden size of 3584 and a total of 32 layers.
@@ 3.3.2 Post-training
After the pre-training phase, we apply supervised fine-tuning (SFT), followed by reinforcement learning from human feedback (RLHF), to further improve the aesthetic quality, motion consistency and structural coherence of the outputs. The post-training process is carried out independently for the text-to-video and image-to-video tasks.
@@
However, prolonged SFT training can lead to rapid overfitting, resulting in degraded prompt-following ability and reduced motion quality. As illustrated in Figure 14, the model’s ability to follow prompts deteriorates during the SFT stage.
@@ 4.1 Quantitative Analysis of Video Generation
In the text-to-video task, Seaweed-7B ranks among the top 2 models in the Elo comparison. It follows the top-ranked model, Veo 2, and outperforms Wan 2.1-14B and Kling 1.6 (HD). Figure 19 presents a detailed comparison with the two leading models, Veo 2.0 and Wan 2.1. These results demonstrate that the Seaweed model, trained with 665,000 H100 GPU hours, achieves competitive performance compared to larger models trained with significantly greater computational resources.
@@ 7 Conclusion
Below are the researchers (individuals who worked full-time on the project) and contributors (individuals who provided assistance to the project). * indicates that the list is ordered alphabetically. The Points of Contact (POCs) in the infrastructure work stream are marked with $\dagger$.
@@
Model
  • Code & Engineering: Shanchuan Lin, Peihao Zhu, Qi Zhao.
  • Modeling Call: Ceyuan Yang.
  • VAE: Yang Zhao, Hao Chen.
  • DiT: Zhijie Lin, Fei Xiao.
  • Post-training: Feng Cheng, Haoyuan Guo, Meng Wei, Zhiwu Qing.
Data
  • Processing Algorithms: Fangyuan Kong, Jiangqiao Yan, Liangke Gui, Lu Qi, Sen Wang, Tuyen Hoang, Ziyan Yang.
  • Processing Pipeline: Zhibei Ma, Sheng Bi.
  • Acquisition: Feilong Zuo, Siyu Zhang.
  • Evaluation & Annotation: Xuejiao Zeng.
@@
Contributors: Bingchuan Li, Chao Liang, Deyao Zhu, Gaojie Lin, Gen Li, Hao He, Jianwen Jiang, Jianyi Wang, Jiaqi Yang, Jiawei Liu, Junfei Xiao, Lijie Liu, Lizhen Wang, Longhao Zhang, Qian He, Ruiqi Xia, Siyu Zhou, Tianshu Hu, Tianxiang Ma, Weilin Huang, Xiaobin Zhuang, Xiaohui Shen, Xinglong Wu, Yongming Zhu, Yuping Wang, Yuwei Guo, Yuxuan Luo, Yuxuan Wang, Zerong Zheng, Zhengkun Rong, Zhuo Chen, Zhuowei Chen.